# Gaussian Process Priors for Boundary Value Problems of Linear Partial Differential Equations

### v1

@@ Abstract
Solving systems of partial differential equations (PDEs) is a fundamental task in computational science, traditionally addressed by numerical solvers. Recent advancements have introduced neural operators and physics-informed neural networks (PINNs) to tackle PDEs, achieving reduced computational costs at the expense of solution quality and accuracy. Gaussian processes (GPs) have also been applied to linear PDEs, with the advantage of always yielding precise solutions. In this work, we propose Boundary Ehrenpreis-Palamodov Gaussian Processes (B-EPGPs), a novel framework for constructing GP priors that satisfy both general systems of linear PDEs with constant coefficients and linear boundary conditions. We explicitly construct GP priors for representative PDE systems with practical boundary conditions. Formal proofs of correctness are provided and empirical results demonstrating significant accuracy improvements over state-of-the-art neural operator approaches.
Gaussian Processes, Nonlinear algebra, Partial Differential Equations
@@ 1 Introduction
Figure 1: Consider the two-dimensional wave equation $\partial_{t}^{2}u-\partial_{x}^{2}u-\partial_{y}^{2}u=0$ with zero boundary condition on a spatial rectangle. This figure shows a sample from our Gaussian process construction B-EPGP. All samples of this Gaussian process analytically satisfy the differential equation, the boundary conditions, and can be conditioned on any finite set of data.
(Systems of) partial differential equations (PDEs) were classically solved using numerical solvers. To be able to find the solution, it needed to be unique. Hence, the PDE system needed enough initial conditions or boundary conditions to ensure this uniqueness. Some techniques for PDEs in machine learning mirror this general approach. For example, neural operators are neural networks trained to solve systems of evolution equations with initial conditions, where the neural network maps a state of the system at time $t$ to time $t+1$. While such approaches require only a fraction of computation time, they do not yet reach the quality of numerical solvers.
A second category of research in machine learning disposes of the uniqueness and instead aims for finding a “good” solution of a PDE system that approximates some data points. An examples are physics informed neural networks (PINN) (Raissi et al., 2019), which are used as neural network models for regression with a usual regression loss and additionally a loss that the regression model satisfies the PDE system at randomly sampled locations in their domain. Also, many Gaussian processes (GPs) based models for linear PDE systems fall in this category.
GPs (Rasmussen & Williams, 2006) are the default functional prior, serve as robust regression models in the context of few data points, and provide calibrated uncertainties of predictions. The bilinear structure of the covariance of GPs is able to encode linear mathematical information in a prior, for example derivatives (Swain et al., 2016; Harrington et al., 2016). Encoding derivatives allowed to construct GPs with realizations in the solution set a given system of linear PDEs with constant coefficients (Lange-Hegermann, 2018; Jidling et al., 2017; Macêdo & Castro, 2008; Scheuerer & Schlather, 2012; Wahlström et al., 2013; Solin et al., 2018; Jidling et al., 2018; Särkkä, 2011), as long as these systems are controllable, i.e. allow for potentials. This construction is non-trivial and makes use fo Gröbner bases for multivariate polynomial rings. The restriction to such controllable systems was first lifted for systems of ordinary differential equations (ODEs) (Besginow & Lange-Hegermann, 2022) and then for PDEs (Harkonen et al., 2023). All of these approaches allow machine learning for a huge and very relevant class of differential equations. Their underlying symbolic nature is not merely physics informed, but all regression functions and samples are exact solutions of the PDE system, making them physics constrained. Hence, such methods improve upon the accuracy of PINN methods by several orders of magnitude.
In practice, PDE systems come with both data and (initial or) boundary conditions. While controllable systems allow the inclusion of boundary conditions (Lange-Hegermann, 2021; Lange-Hegermann & Robertz, 2022) into GP approaches, at the cost of using Gröbner bases over Weyl algebra, this is not possible for the general case of linear PDE systems with constant coefficients. This paper develops an algorithmic construction of GP priors inside the solution set of any given system of (ordinary or partial) linear differential equations with constant coefficients with linear boundary conditions, which we call B-EPGP for Boundary Ehrenpreis-Paralamodov Gaussian Processes. We built upon the (infinite dimensional) basis of the solution space suggested by the Ehrenpreis-Palamodov fundamental principle and map this basis to a basis of solutions of the PDE which satisfy the boundary condition.
The major contributions of this paper are as follows:
@@
    We give a general construction of GP priors for solutions of systems of linear PDEs with constant coefficients, including boundary conditions.
@@
    We explicitly construct the priors for several PDE systems with suitable boundary conditions.
@@
    We prove correctness and convergence properties of our construction. In particular, all our GP realizations satisfy the system of PDEs and boundary conditions.
  4. 4.
    We empirically demonstrate the superiority of our approach over state of the art neural operator approaches.
For the relevant code and videos, we refer to the github repository or the ancillary files of the arXiv version.
2 Gaussian Process Priors from the Ehrenpreis-Palamodov Theorem
2.1 Ehrenpreis-Palamodov Theorem
An important difference between linear and nonlinear ODE is that solutions to linear ODE systems with constant coefficients are linear combinations of exponential-polynomial solutions. If one looks at the ODE $y^{\prime\prime\prime}-3y^{\prime}+2y=0$, its exponential-polynomial solutions are of the form $e^{x},\,xe^{x},\,e^{-2x}$, where the “frequencies” $1$ and $-2$ are the roots of the characteristic polynomial $z^{3}-3z+2=(z-1)^{2}(z+2)$.
Remarkably, this generalizes to systems of linear partial differential equations (PDE), as follows from the Ehrenpreis–Palamodov Fundamental Principle. Take the example of the 1-D heat equation, $\partial_{1}u-\partial_{2}^{2}u=0$. Its exponential-polynomial solutions will be of the form $e^{x_{1}z_{1}+x_{2}z_{2}}$, where $z_{i}\in\mathbb{C}$ satisfy the characteristic equation $z_{1}^{1}-z_{2}^{2}=0$. This equations is obtained by replacing the partial derivative $\partial_{i}$ with the polynomial variable $z_{i}$, a form of Fourier transform. We say that $z$ lies in the characteristic variety $V=\{(z_{2}^{2},z_{2})\colon z_{2}\in\mathbb{C}\}$. For simplicity, we state a particular case of the Ehrenpreis–Palamodov theorem.
@@ 2.1 Ehrenpreis-Palamodov Theorem
Let $\Omega\subset\mathbb{R}^{n}$ be an open convex set. Let $A\in\mathbb{C}[\partial_{1},\ldots,\partial_{n}]$ and $V=\{z\in\mathbb{C}^{n}\colon A(z)=0\}$. Suppose that $V$ is an irreducible complex variety with no multiplicity. Then the linear span of $\{e^{x\cdot z}\}_{z\in V}$ is dense in the space of smooth solutions of $A(\partial)u(x)=0$ in $\Omega$.
We will present this result in full generality in Appendix C, which accounts for systems, reducible varieties and multiplicities. We clarify that the solutions we construct here are $\ker_{C^{\infty}(\Omega)}A=\{u\in C^{\infty}(\Omega)\colon A(\partial)u=0\}$ with its classic Fréchet topology. For the remainder of this section, we keep the assumptions of Theorem 2.1.
2.2 (Sparse) Ehrenpreis-Palamodov Gaussian Processes
(Harkonen et al., 2023) defined a Gaussian Process (GP) prior with realizations of the form $f(x)=\sum_{j=1}^{r}w_{j}e^{x\cdot z_{j}}$ where all “frequencies” $z_{j}$ lie in the complex variety $V$ and $w_{j}\sim\mathcal{N}(0,\sigma_{j}^{2})$. This required careful local parametrization of $V$ to enable use of SGD to optimize the $z_{j}$ and $\sigma_{j}^{2}$. We dubbed the method EPGP111(Harkonen et al., 2023) coined two terms, EPGP and S-EPGP for two different implementations of our method; here we will refer to either algorithm simply by EPGP. and used it very successfully to reconstruct solutions from data at points, in some cases obtaining results of several orders of magnitude better than state of the art PINN methods.
@@ 3 EPGP for Boundary Value Problems
As mentioned above, EPGP excels at reconstructing solutions from given data. This has worked very well for fitting solutions to 1-D heat equation in from only 16 data points (Harkonen et al., 2023). We also obtained very accurate results in physically meaningful situations when we fitted the solutions from given initial conditions. This was the case for several examples involving the 2-D wave equation, $\partial_{t}^{2}u-\partial_{x}^{2}u-\partial_{y}^{2}u=0$.
In order to tackle realistic problems, boundary conditions must also be included in EPGP.
3.1 Naive baseline: model boundaries using data
It is also possible to incorporate boundary conditions in EPGP in the following manner. Consider for example the initial-boundary value problem for the 1-D heat equation,
@@ 3.1 Naive baseline: model boundaries using data
A direct implementation of EPGP would require setting
$u(t,x)=\sum_{j=1}^{r}w_{j}e^{z_{j}^{2}t+z_{j}x}$
for $z_{j}\in\mathbb{C}$ and imposing the data $u(0,X)=f(X)$ as well as $u(T,\pm 2)=0$ for finite but large enough sets $X\subset[-2,2]$ and $T\subset[0,1]$. See Section F.3 for an implementation of this method for the 2-D wave equation.
This method works reasonably well in a low dimensional setting, when the boundary can be well-described by few data points due to its low volume. In higher dimension, the curse of dimensionality prevents adding enough data points to ensure suitable accuracy. Hence, this baseline falls short when compared with our new method, which imposes the boundary condition directly on the basis elements.
@@ 3.2 B-EPGP: Bases satisfying boundary conditions in halfspaces
In the first instance, we work with the simplest boundaries, i.e., halfspaces. In fact, by a change of variable, we can choose $\Omega=\operatorname*{\mathbb{R}}^{n}_{+}:=\mathbb{R}^{n-1}\times[0,\infty)$. We will later generalize this approach to more general boundaries and show examples in these general cases.
We construct GP priors with realizations which satisfy
@@
where $B\in\mathbb{C}[\partial]^{h\times 1}$ is a linear PDE operator which represents the boundary conditions. Starting with EPGP (cf. Theorem 2.1) any realization
$f(x)=\sum_{j=1}^{r}w_{j}e^{x\cdot z_{j}}$
in the set of solutions of (1) should also satisfy the boundary condition exactly. In contrast, the naive baseline from Section 3.1 satisfies the boundary condition approximately.
Write $x=(x^{\prime},x_{n})$ where $x^{\prime}\in\operatorname*{\mathbb{R}}^{n-1}$, $x_{n}\in\operatorname*{\mathbb{R}}$. Setting $x_{n}=0$ in $B(z)f=0$ leads to the algebraic condition
$\sum_{j=1}^{r}w_{j}B(z_{j})e^{x^{\prime}\cdot z^{\prime}_{j}}=0\text{ for all }x^{\prime}\in{\mathbb{R}^{n-1}},$
where $z_{j}^{\prime}$ denotes the first $n-1$ components of $z_{j}$. For this to hold, we must have that $z_{j}^{\prime}$ is constant with respect to $j$. Hence, $\sum_{j=1}^{r}w_{j}B(z_{j})=0$ for all $x^{\prime}\in{\mathbb{R}^{n-1}}$. This enables us to write an algorithm to construct basis elements for solutions of (1).
Let $z^{\prime}\in\mathbb{C}^{n-1}$ be such that there exists $z_{n}\in\mathbb{C}$ for which $(z^{\prime},z_{n})\in V$. Let $V^{\prime}$ be the set of such $z^{\prime}$, i.e., the projection of the characteristic variety $V$ on $\mathbb{C}^{n-1}$. Denote by $S_{z^{\prime}}=\{z\in V\colon z=(z^{\prime},z_{n})\text{ for some }z_{n}\in\mathbb{C}\}$ the non-empty fibers of this projection, which can be computed by non-trivial Gröbner basis methods (Barakat & Lange-Hegermann, 2022). Then find $w_{z}\in\mathbb{C}$ such that
$\sum_{z\in S_{z^{\prime}}}w_{z}B(z)=0.$ (2)
The set of possible $w_{z}$ are the left syzygies when vertically stacking the matrices $B(z)$ for $z\in S_{z^{\prime}}$. These syzygies are computable by standard Gröbner basis algorithms.
We obtain the basis
$\left\{\sum_{z\in S_{z^{\prime}}}w_{z}B(z)e^{x\cdot z}\colon\sum_{z\in S_{z^{\prime}}}w_{z}B(z)=0\right\}_{z^{\prime}\in V^{\prime}}.$ (3)
The basis (3) can be computed explicitly by hand for the heat and wave equations with either Dirichlet or Neumann boundary conditions.
@@
If we set $B=1$, we will obtain $\sum w_{z}=0$ in (2). This is a broadly used condition with which we can solve explicitly. ∎
@@
Consider the 1-D heat equation $\partial_{t}u-\partial_{x}^{2}u=0$ with Dirichlet boundary condition, $B=1$ and $u(t,0)=0$. Start with $z=(\zeta^{2},\zeta)\in V$ and note that $z^{\prime}=\zeta^{2}$, so that $S_{z^{\prime}}=\{\pm\zeta\}$. We conclude that in this case the basis elements are $\{e^{t\zeta^{2}+x\zeta}-e^{t\zeta^{2}-x\zeta}\}_{\zeta\in\mathbb{C}}.$
@@
Similarly, for the 2-D wave equation $\partial_{t}^{2}u-\partial_{x}^{2}u-\partial_{y}^{2}u=0$ with Dirichlet boundary condition $u(t,0,y)=0$, we obtain the basis elements $\{e^{\pm\sqrt{a^{2}+b^{2}}t+ax+by}-e^{\pm\sqrt{a^{2}+b^{2}}t-ax+by}\}_{a,b\in\mathbb{C}}$. ∎
@@
Another widely used boundary condition is Neumann, i.e., $B(z)=z_{n}$, which leads to $\sum w_{z}z_{n}=0$. Comparing to the previous examples, this leads to the basis $\{e^{t\zeta^{2}+x\zeta}+e^{t\zeta^{2}-x\zeta}\}_{\zeta\in\mathbb{C}}$ for 1-D heat equation (Example 3.2) and $\{e^{\sqrt{a^{2}+b^{2}}t+ax\pm by}+e^{\sqrt{a^{2}+b^{2}}t+ax\mp by}\}_{a,b\in\mathbb{C}}$ for 2-D wave equation (Example 3.3). ∎
These calculations can be performed in arbitrary dimensions and for arbitrary hyperplanes, see Appendix A. This basis replaces the basis of exponential functions in EPGP. Importantly, we can prove that our method constructs approximations of all solutions to the boundary value problem:
@@
We will prove this in Appendix B, see Appendix C for details on the topology with respect to which we have density. Relevant computational results are in Sections 4 and 5.1.
This method has substantial advantages over the direct method in Section 3.1. First, we construct solutions which satisfy the boundary condition exactly, not approximately; this improves precision and reduces the risk of overfitting. Second, no additional data points are necessary to model the boundary conditions; this improves computation time. Our numerical experiments demonstrate this clearly.
3.3 B-EPGP with linear boundaries
We extend B-EPGP to more general polygonal boundaries. Consider the following computational steps to generate a basis for the solutions set given boundary conditions with multiple halfspaces with boundaries $H_{k}$.
[…]
Step 2. IF all elements of $b(z)$ satisfy the boundary condition, THEN stop. ELSE repeat Step 2 for each $\zeta\in\varphi(z)$.
These computational steps do not terminate in general. However, if the reflection group generated by the $H_{k}$ is finite, this process obviously terminates. An important case where this reflection group is finite is the case of all boundaries being orthogonal to each other. We consider cases of infinite reflection groups below in Example 3.7, where we can still construct a basis. In all remaining examples below, these computational steps terminate.

### v2

@@ Abstract
Working with systems of partial differential equations (PDEs) is a fundamental task in computational science. Well-posed systems are addressed by numerical solvers or neural operators, whereas systems described by data are often addressed by PINNs or Gaussian processes. In this work, we propose Boundary Ehrenpreis–Palamodov Gaussian Processes (B-EPGPs), a novel probabilistic framework for constructing GP priors that satisfy both general systems of linear PDEs with constant coefficients and linear boundary conditions and can be conditioned on a finite data set. We explicitly construct GP priors for representative PDE systems with practical boundary conditions. Formal proofs of correctness are provided and empirical results demonstrating significant accuracy and computational resource improvements over state-of-the-art approaches.
@@ 1 Introduction
Classically, systems of partial differential equations (PDEs) have been solved using numerical methods, which require the solution to be uniquely determined. This uniqueness is typically ensured by prescribing a sufficient number of initial or boundary conditions. Neural operators adopt a similar strategy, and learn the evolution of PDE solutions, mapping the system state at time $t$ to the state at time $t+1$, given appropriate initial conditions. While such methods significantly reduce computational cost, their accuracy falls short of that achieved by traditional numerical solvers.
A second line of research in machine learning abandons the requirement of uniqueness and instead seeks a “good” approximate solution that fits observed data, i.e. regression inside of the solution set of PDEs. This paper falls in this line of work. A prominent example is physics-informed neural networks (PINNs) (raissi2019physics, ), which combine standard regression loss with an additional penalty for violating the PDE at sampled points in the domain. Many Gaussian process (GP) models for linear PDEs also follow this approach.
GPs (RW, ) are a standard choice for functional priors, offering robust regression with few data points and calibrated uncertainty estimates. Their bilinear covariance structure can encode linear properties, such as derivatives (swain2016inferring, ; harrington2016differential, ), enabling the construction of GPs whose realizations lie in the solution set of linear PDEs with constant coefficients (LH_AlgorithmicLinearlyConstrainedGaussianProcesses, ; LinearlyConstrainedGP, ; MacedoCastro2008, ; scheuerer2012covariance, ; Wahlstrom13modelingmagnetic, ; MagneticFieldGP, ; jidling2018probabilistic, ; sarkka2011linear, ), provided the system is controllable, i.e., admits potentials. This construction, relying on Gröbner bases for multivariate polynomial rings, was extended beyond controllable systems (harkonen2023gaussian, ). These methods yield machine learning models whose predictions are not merely physics-informed, but physics constrained: all regression functions are exact solutions of the PDE system. As a result, they surpass PINNs in accuracy by several orders of magnitude.
@@
Figure 1: Consider the 2D wave equation $\partial_{t}^{2}u-\partial_{x}^{2}u-\partial_{y}^{2}u=0$ with zero boundary conditions on a rectangular spatial domain. This figure shows a sample drawn from our B-EPGP construction. Every sample analytically satisfies both the differential equation and the boundary conditions, and can be conditioned on any finite set of observations.
In practice, PDE systems are typically accompanied by both data and (initial or) boundary conditions. For controllable systems, such boundary conditions can be incorporated into GP models using Gröbner bases over the Weyl algebra (LH_AlgorithmicLinearlyConstrainedGaussianProcessesBoundaryConditions, ). However, this approach does not extend to general linear PDEs with constant coefficients. In this paper, we introduce an algorithmic construction of GP priors—called Boundary Ehrenpreis–Palamodov Gaussian Processes (B-EPGPs)—that reside within the solution space of any such system, including linear boundary conditions. B-EPGP builds on the infinite-dimensional basis provided by the Ehrenpreis–Palamodov theorem and transforms it into a basis of solutions that also satisfy the boundary conditions. Our main contributions are:
@@
    We present a general algorithmic construction of a novel model class: GP priors for solutions of linear PDE systems with constant coefficients and linear boundary conditions.
@@
    We provide explicit constructions of such priors for several representative PDEs with practical boundary setups.
@@
    We thoroughly prove correctness and convergence of our approach, in particular that these GP priors have realizations that are dense in the PDE solution.
While our approach is mainly a regression model, we empirically demonstrate that it significantly outperforms state-of-the-art neural operator models by orders of magnitude in accuracy and computation time. Our approach scales mildly with the dimension, in particular we can handle the 3D wave equation (with 4 inputs) with minimal boundary data, avoiding the exponential boundary discretization cost of traditional solvers.
2 Gaussian Process Priors from the Ehrenpreis–Palamodov Theorem
A key distinction between linear and nonlinear ODEs is that solutions to linear systems with constant coefficients are linear combinations of exponential-polynomial functions. For example, the ODE $y^{\prime\prime\prime}-3y^{\prime}+2y=0$ has linear combinations of $e^{x}$, $xe^{x}$, and $e^{-2x}$ as solutions, where the “frequencies” $1$ and 2-2 are the roots of the characteristic polynomial $z^{3}-3z+2=(z-1)^{2}(z+2)$.
Remarkably, this structure extends to systems of linear PDEs with constant coefficients. Consider the 1D heat equation $\partial_{1}u-\partial_{2}^{2}u=0$. Its exponential-polynomial solutions take the form $e^{x_{1}z_{1}+x_{2}z_{2}}$, where the frequency vector $(z_{1},z_{2})$ satisfies the characteristic equation $z_{1}-z_{2}^{2}=0$. This equation arises substituting each derivative $\partial_{i}$ with a complex variable $z_{i}$, akin to a Fourier transform. The set of such $z$ defines the characteristic variety, here given by $V=\{(z_{2}^{2},z_{2})\colon z_{2}\in\mathbb{C}\}$. We state the special case of the Ehrenpreis–Palamodov theorem for irreducible characteristic varieties without multiplicities. We refer to Appendix E for the general case, which involves linear combinations over the components of the variety and correction terms (as the factor $x$ in $xe^{x}$ above) for multiplicities.
@@ 2 Gaussian Process Priors from the Ehrenpreis–Palamodov Theorem
Let $\Omega\subset\mathbb{R}^{n}$ be an open convex set. Let $A\in\mathbb{C}[\partial_{1},\ldots,\partial_{n}]$ and define the characteristic variety $V=\{z\in\mathbb{C}^{n}\colon A(z)=0\}$. Suppose $V$ is irreducible and has no multiplicities. Then, the linear span of $\{e^{x\cdot z}\}_{z\in V}$ is dense in the space of smooth solutions of $A(\partial)u(x)=0$ on $\Omega$.
The solutions we construct are $\ker_{C^{\infty}(\Omega)}A=\{u\in C^{\infty}(\Omega)\colon A(\partial)u=0\}$ endowed with the classic Fréchet topology. Until Appendix E, we assume the simplifying conditions of Theorem 2.1 hold.
(harkonen2023gaussian, ) defined a Gaussian Process (GP) prior with realizations of the form $f(x)=\sum_{j=1}^{r}w_{j}e^{x\cdot z_{j}}$ where all “frequencies” $z_{j}$ lie in the complex variety $V$ and $w_{j}\sim\mathcal{N}(0,\sigma_{j}^{2})$. Local parametrization of $V$ enabled the use of SGD to optimize the $z_{j}$ and $\sigma_{j}^{2}$. The method was dubbed EPGP111(harkonen2023gaussian, ) coined two terms, EPGP and S-EPGP; we will refer to either algorithm by EPGP. and successfully build probability distributions of PDE solutions from data at points, in some cases obtaining results of several orders of magnitude better than state-of-the-art PINN methods.
This paper follows this general idea of a probabilistic model for solutions of PDE systems and replaces the basis functions $e^{x\cdot z_{j}}$ from EPGP by new basis functions that satisfy boundary conditions.
@@ 3 EPGP for Boundary Value Problems
We introduce and demonstrate B-EPGP on realistic problems with boundary conditions. Additional concrete examples can be found in Appendices H.1H.2H.3J.1J.3KL.2,L.1.
Remark 3.1.
Boundary conditions can be incorporated into EPGP by conditioning the GP on data at the boundary. Consider, for example, the initial-boundary value problem for the 1D heat equation:
@@
A direct EPGP implementation sets $u(t,x)=\sum_{j=1}^{r}w_{j}e^{z_{j}^{2}t+z_{j}x}$ with $z_{j}\in\mathbb{C}$, and fits the initial and boundary data at finite sets $X\subset[-2,2]$ and $T\subset[0,1]$. We apply this approach to the 2D wave equation in Section H.3. This method performs reasonably well in low dimensions, where boundaries can be approximated with few data points. However, in higher dimensions, the curse of dimensionality makes this approach highly inefficient, as $g^{d-1}$ points are needed for a boundary in $d$-dimensional space with a grid containing $g$ points in each dimension. This corresponds to the typical curse of dimensionality in numerical approximations. B-EPGP encodes the boundary condition directly into the basis elements, avoiding the curse of dimensionality completely when encoding the boundary. B-EPGP is not affected by the curse of dimensionality even for inhomogeneous boundary conditions, see Appendix L.
@@ 3.1 B-EPGP: Bases satisfying boundary conditions in halfspaces
We begin our introduction of B-EPGP with the simplest class of boundaries: halfspaces. By a change of variables, we may assume the domain is $\Omega=\mathbb{R}^{n}_{+}:=\mathbb{R}^{n-1}\times[0,\infty)$. We will later generalize this approach to more complex boundaries and provide corresponding examples.
Our goal is to construct GP priors whose realizations satisfy
@@
where $B\in\mathbb{C}[\partial]^{h\times 1}$ is a linear PDE operator representing the boundary condition. Starting from Theorem 2.1, we consider realizations of the form $f(x)=\sum_{j=1}^{r}w_{j}e^{x\cdot z_{j}},$ which we aim to constrain to satisfy both the PDE and the boundary condition exactly. In contrast, the baseline approach from Remark 3.1 enforces the boundary condition only approximately and at a much higher cost. We stress that we are interested in underdetermined systems (1), i.e., systems with many solutions. This means that our method constructs probability distributions on infinite dimensional spaces – the kernels (1).
For $x=(x^{\prime},x_{n})\in\operatorname*{\mathbb{R}}^{n-1}\times\operatorname*{\mathbb{R}}$, evaluating $B(\partial)f=0$ at $x_{n}=0$ yields $\sum_{j=1}^{r}w_{j}B(z_{j})e^{x^{\prime}\cdot z^{\prime}_{j}}=0\text{ for all }x^{\prime}\in{\mathbb{R}^{n-1}},$ where $z_{j}^{\prime}$ denotes the first n1n-1 components of $z_{j}$. For this identity to hold, we must have that $z_{j}^{\prime}$ is constant with respect to $j$. Hence, $\sum_{j=1}^{r}w_{j}B(z_{j})=0$ for all $x^{\prime}\in{\mathbb{R}^{n-1}}$. This enables us to write an algorithm to construct basis elements for solutions of (1).
Let $z^{\prime}\in\mathbb{C}^{n-1}$ be such that there exists $z_{n}\in\mathbb{C}$ with $(z^{\prime},z_{n})\in V$, and define $V^{\prime}$ as the projection of the characteristic variety $V$ onto $\mathbb{C}^{n-1}$. For each $z^{\prime}\in V^{\prime}$, let $S_{z^{\prime}}=\{z\in V\colon z=(z^{\prime},z_{n})\text{ for some }z_{n}\in\mathbb{C}\}$ denote the non-empty fiber over $z^{\prime}$. These fibers can be computed using Gröbner basis with negligible complexity in practice, see Appendix B. To ensure the boundary condition is satisfied, we seek weights $w_{z}\in\mathbb{C}$ such that
$\sum\nolimits_{z\in S_{z^{\prime}}}w_{z}B(z)=0.$ (2)
The admissible choices of $w_{z}$ form the left syzygy module of the vertically stacked matrices $B(z)$ for $z\in S_{z^{\prime}}$, which is again computable via Gröbner basis algorithms with negligible complexity, see again Appendix B. For each $z^{\prime}\in V^{\prime}$, we obtain the basis
$\left\{\sum\nolimits_{z\in S_{z^{\prime}}}w_{z}B(z)e^{x\cdot z}\colon\sum\nolimits_{z\in S_{z^{\prime}}}w_{z}B(z)=0\right\}_{z^{\prime}\in V^{\prime}}$ (3)
of terms for the frequency $z^{\prime}$. We use a finite union of such basis vectors to perform linear regression as in EPGP. The bases (3) can be computed explicitly by hand for the heat and wave equations with either Dirichlet or Neumann boundary conditions.
@@
If we set $B=1$, we obtain $\sum w_{z}=0$ in (2). This is a broadly used condition with which we can solve explicitly. ∎
@@
Consider the 1D heat equation $\partial_{t}u-\partial_{x}^{2}u=0$ with Dirichlet boundary condition, $B=1$ and $u(t,0)=0$. Start with $z=(\zeta^{2},\zeta)\in V$ and note that $z^{\prime}=\zeta^{2}$, so that $S_{z^{\prime}}=\{\pm\zeta\}$. We conclude that in this case the basis elements are $\{e^{t\zeta^{2}+x\zeta}-e^{t\zeta^{2}-x\zeta}\}_{\zeta\in\mathbb{C}}.$
@@
Similarly, for the 2D wave equation $\partial_{t}^{2}u-\partial_{x}^{2}u-\partial_{y}^{2}u=0$ with Dirichlet boundary condition $u(t,0,y)=0$, we obtain the basis $\{e^{\pm\sqrt{a^{2}+b^{2}}t+ax+by}-e^{\pm\sqrt{a^{2}+b^{2}}t-ax+by}\}_{a,b\in\mathbb{C}}$. ∎
@@
Another widely used boundary condition is Neumann, i.e., $B(z)=z_{n}$, which leads to $\sum w_{z}z_{n}=0$. Comparing to the previous examples, this leads to the basis $\{e^{t\zeta^{2}+x\zeta}+e^{t\zeta^{2}-x\zeta}\}_{\zeta\in\mathbb{C}}$ for 1D heat equation (Example 3.3) and $\{e^{\sqrt{a^{2}+b^{2}}t+ax\pm by}+e^{\sqrt{a^{2}+b^{2}}t+ax\mp by}\}_{a,b\in\mathbb{C}}$ for 2D wave equation (Example 3.4). ∎
These calculations can be performed in arbitrary dimensions and for arbitrary hyperplanes, see Appendix C. This basis replaces the basis of exponential functions in EPGP. Importantly, we can prove that our method constructs approximations of all solutions to the boundary value problem:
@@
We prove this in Appendix D, see Appendix E for details on the topology with respect to which we have density. Relevant computational results are in Sections 5 and Appendices HI.
3.2 B-EPGP with polygonal boundaries
B-EPGP handles polygonal boundary conditions given by $k$ halfspaces $H_{i}$ for $1\leq i\leq k$ as follows. We rephrase the computational steps from Section 3.1 in two parts, starting from a single EPGP basis function $e^{\widehat{z}\cdot x}$ with frequency $\widehat{z}\in V$:
  • (i)
    Use the geometric arguments from above to collect all relevant frequency vectors in the fiber $S_{\widehat{z}^{\prime}}\ni\widehat{z}$ in the context of the single halfspace.
  • (ii)
    Determine all linear combinations of the basis functions $e^{z\cdot x}$ for $z\in S_{\widehat{z}^{\prime}}$ that satisfy the boundary condition, using the syzygy computation explained above.
To extend this to polygonal boundaries, we refine step (i) as follows, while keeping step (ii) unchanged:
  • (i’)
    Initialize a singleton frequency set $S=\{\widehat{z}\}$. Iteratively apply the construction from (i) to $z\in S$ and a halfspace $H_{i}$, and add any resulting frequencies to $S$. Repeat until $S$ stabilizes, i.e., no new frequencies are introduced for any combination $z\in S$ and $H_{i}$, $1\leq i\leq k$.
These steps in (i’) do not terminate in general. We demonstrate how to still construct a basis with non-termination in Example 3.8. In all remaining examples below, the computation terminates.